OpenStudy (anonymous):

Factor and solve 2u^2 + 9u = 18 for u.

OpenStudy (anonymous):

$$2u^2+9u-18=0\\2u^2+12u-3u-18=0\\2u(u+6)-3(u+6)=0\\(2u-3)(u+6)=0$$so since our product is \(0\) it follows at least one factor must be \(0\):$$2u-3=0\\2u=3\\u=\frac32$$and$$u+6=0\\u=-6$$

OpenStudy (anonymous):

Thank you!